OpenStudy (anonymous):

Differentiate the function y = f(x) = x+1 divided by x−1. there are no parentheses if it matters. my answer was -2 divided by x squared - 2x + 1. ( pretty sure i missed it)

OpenStudy (anonymous):

f(x) = u/v u = (x+1) v = (x-1) maybe you used the formula : f'(x) = (u'v - v'u)/(v^2) so lets calculate it : u' = 1 v' = 1 so u'v = 1*(x-1) v'u = 1*(x+1) v^2 = (x-1)^2 u'v - v'u = x-1 -(x+1) = -2 put all together : f'(x) = -2/((x-1)^2)

OpenStudy (anonymous):

so i was correct.

OpenStudy (anonymous):

yes :) now ive noticed lol i could spare time :|